# SafeVLA: Towards Safety Alignment of Vision-Language-Action Model via Safe Reinforcement Learning

### v1

@@ 5.1 Experiment Setup
Perturbation SR $\uparrow$ RC $\downarrow$ OC $\downarrow$ SR $\uparrow$ RC $\downarrow$ OC $\downarrow$ SR $\uparrow$ RC $\downarrow$ OC $\downarrow$
SafeVLA 0.865 1.698 0.156 0.915 1.042 0.154 0.637 7.93 1.054
+Color 0.804 2.864 0.231 0.902 1.634 0.182 0.602 14.5 0.837
+Light 0.833 2.172 0.318 0.928 0.562 0.125 0.605 7.433 1.083
+Material 0.839 2.798 0.185 0.916 0.482 0.156 0.653 7.077 1.167
+All 0.817 2.923 0.289 0.903 0.267 0.139 0.589 11.578 0.918
Average -0.042 +0.991 +0.099 -0.027 -0.306 -0.004 -0.025 +2.217 -0.053
Baselines
We conducted comparative evaluations by analyzing three categories of methods: 1) a combination of imitation learning (IL) and reinforcement learning (RL), 2) pure IL, and 3) standalone RL. The categories of each method are summarized in the first two columns of Table 1. The SPOC model series (Ehsani et al., 2024), which serves as the VLA foundation model trained through large-scale imitation learning and forms the base model for our SafeVLA experiments; FLaRe (Hu et al., 2024), a reinforcement learning fine-tuning approach using foundation models, with its reward-shaping variant FLaRe-RS directly incorporating cost as a penalty in the reward function; and Poliformer (Zeng et al., 2024), a navigation-specific RL method developed via end-to-end training without relying on pre-existing priors.
5.2 Main Results
As shown in Table 1 and Table 2 that SafeVLA achieves the highest task performance and the lowest cumulative cost across all tasks, outperforming existing state-of-the-art methods in both safety and task performance with average improvements of 83.58% and 3.85%, respectively. Furthermore, the improvements in safety and task performance achieved through SafeVLA alignment are maintained even under highly OOD scenarios. From a qualitative analysis perspective, we observe in Figure 2 that SafeVLA decouples the optimization objectives of safety and task performance by treating safety as an independent dimension. This decoupling ensures that, even in instances of task failure, a safe behavioral paradigm is preserved, thereby eliminating the long-tail distribution of safety risks.
5.2.1 Qualitative Analysis
In Figure 2, we present the cumulative cost distributions of the robot trajectories generated on the test sets after fine-tuning the models using SafeVLA and FLaRe (which optimizes only task performance). The results show that SafeVLA significantly improves the cost distribution, reducing the probability of high-risk unsafe behaviors (cumulative cost >10) to zero. It also lowers the upper bound of unsafe behaviors to $1/35$ of that seen in the FLaRe model, significantly reducing the upper bound of risk while greatly enhancing overall safety.
Additionally, compared to approaches that optimize solely for task performance, SafeVLA decouples and concurrently optimizes both safety and task performance objectives. This capability is evidenced by the strong correlation between the cumulative cost distribution and task success observed in the FLaRe model, as shown in Figure 2. In successful cases, the severity and frequency of unsafe behaviors are minimal, whereas in failure cases, unsafe behaviors are more uniformly distributed. For the SafeVLA model, however, the cumulative cost distribution of robot trajectories remains consistent regardless of task success or failure. This suggests that the model has learned a safe behavior paradigm that is independent of task success, effectively decoupling the safety and task objectives. Analysis of the logistic regression curve for the FLaRe model reveals a clear decline in task success rate as cost increases, and a Pearson correlation T-test at a 1% significance level further confirms that while the SafeVLA model’s cumulative cost is independent of success, this independence is strictly rejected for the FLaRe model (see Appendix A for further details).
5.2.2 Quantitative Evaluation
In Table 1, we compare SafeVLA with alternative methods across three key metrics: success rate (SR), cumulative robot cost (RC), and cumulative object cost (OC). These metrics are critical for evaluating both the safety and the task performance of robot trajectories. On average, SafeVLA has been shown to improve overall safety by 83.58%, as measured by RC and OC, compared to the baseline IL and other RL-based methods. As illustrated in Figure 4, safety has been consistently and substantially enhanced across all tasks and different rooms.
In terms of task performance, SafeVLA clearly outperforms the IL baseline and either matches or exceeds the performance of other RL-based methods. This finding indicates that SafeVLA effectively balances the dual objectives of safety and task performance, in contrast to models that focus only on the latter.
Moreover, as shown in Table 2, robustness is demonstrated by the SafeVLA-trained model when it is subjected to various OOD disturbances, with the original safety and task performance being consistently maintained. This suggests that the benefits brought by SafeVLA are not only effective in controlled environments but also generalize well to challenging and unpredictable conditions.
Table 3: Performance of SafeVLA with different cumulative cost thresholds in Safety-ObjNav. This table shows the success rate (SR), cumulative robot cost (RC), and cumulative object cost (OC) for various threshold settings. The original threshold value of 12.356 was derived from the cumulative cost after the convergence of the FLaRe model, with subsequent thresholds set as percentages of this value.
Threshold SR $\uparrow$ RC $\downarrow$ OC $\downarrow$ Cumulative Cost $\downarrow$
10% 0.822 0.964 0.198 1.162
20% 0.865 1.698 0.156 1.854
50% 0.868 5.5 0.409 5.909
5.3 Impact of Cost Threshold
As shown in equation (7), the hyperparameter $b_{i}$ is used to regulate the strictness of the constraint $c_{i}$. To evaluate the effect of $b_{i}$ on performance, we conducted an ablation study in the Safety-ObjNav environment, setting $b_{i}$ to 50%, 20%, and 10% of the converged cumulative cost in FLaRe, respectively. As reported in Table 3, the cumulative cost remains effectively constrained within the specified range across the various threshold values, while the success rate is maximized. These findings indicate that SafeVLA achieves a robust balance between safety and task performance under differing levels of constraint strictness.
6 Conclusion and Outlooks
In this work, we overcome a fundamental limitation of existing vision-language-action models (VLAs), namely, that even the most advanced models still haven’t made safety an explicit part of their design. To address this gap, we leverage the constrained Markov decision process paradigm, integrating safety directly into the model by contrasting safety-positive and safety-negative samples within simulated environments. To the best of our knowledge, SafeVLA is the first algorithm to explicitly integrate safety modeling into VLAs. This novel integration provides a robust solution for balancing safety and task performance across diverse tasks and out-of-distribution scenarios. As a result, SafeVLA outperforms the state-of-the-art method in both safety and task performance, achieving average improvements of 83.58% and 3.85%, respectively. Additionally, compared to the Reward Shaping baseline, SafeVLA’s explicit safety modeling paradigm facilitates a more effective balance between safety and task performance. This approach ensures that the model maximizes performance while adhering to predefined safety constraints.
Limitations and Future Work
Despite its promising results, this work has several limitations. A primary limitation is that, although we have thoroughly demonstrated the effectiveness of SafeVLA in simulated experiments, we have not yet validated the performance of models fine-tuned on simulated data in real-world environments and robotic platforms due to constraints in available facilities and equipment. Previous research has shown that fine-tuning a model with sufficiently diverse simulated data and then deploying it directly into real-world scenarios is feasible (Ehsani et al., 2024; Hu et al., 2024). This assumption is particularly important in robot safety research, as collecting extensive safety-related negative samples in real-world settings would be prohibitively expensive. Additionally, our current safety constraints are applied continuously throughout the robot’s task execution, rather than being explicitly linked to specific task instructions. To ensure safety in dynamic environments, the problem may need to be reformulated by incorporating constraints that account for both the dynamic nature of the environment and the safety requirements specified through language-based instructions. Moreover, to guarantee the reliability of robot trajectories in real-world settings, external safety measures, such as physical barriers or filtering mechanisms, will be necessary to provide robust protection.
Future work will focus on validating SafeVLA in more complex real-world robot environments. Physical interactions in real-world settings present significant challenges, particularly the sim-to-real gap and the irreversible consequences of even minor failures. We aim to incorporate dynamic safety constraints that adapt to changing conditions and human interactions. To address these challenges, we plan to develop robust uncertainty estimation methods for real-time risk assessment and implement adaptive safety boundaries that respond to environmental dynamics. Furthermore, exploring comprehensive safety mechanisms, including both algorithmic safeguards and physical safety measures, will be crucial to ensuring the robustness of the system in real-world deployments.
Impact statement
@@ Appendix A Additional Empirical Results
In Figure 3, we present the logistic regression analysis of task success probability as a function of cumulative cost for the SafeVLA and FLaRe models, and in Table 4, we provide the correlation coefficients and significance levels for these models.
Figure 3: Logistic regression analysis of task success versus cumulative cost. Left: Logistic regression analysis of task success probability as a function of cumulative cost for the SafeVLA model. The model maintains a relatively high probability of success across different cost levels, indicating its robustness in handling cost variations. Right: Logistic regression analysis of task success probability for the FLaRe baseline model. A sharp decline in success probability is observed as cumulative cost increases, suggesting a stronger correlation between cumulative cost and task failure in the baseline model.
@@
Appendix B Implementation Details and Hyperparameters
B.1 Details of SafeVLA Training
@@
Drawing inspiration from Safe-RLHF (Dai et al., 2023), the learning phase of SafeVLA involves iteratively solving the min-max problem defined in Equation (9). Specifically, we alternate between updating the VLA model parameters, $\theta$, and the Lagrange multipliers, $\lambda$. The cost and reward functions are defined as follows:
$r_{t}$ $=(l,s_{t}),$ (10)
$c_{t}$ $=c_{\text{object}}(s_{t},a_{t},s_{t+1})+c_{\text{robot}}(s_{t},a_{t},s_{t+1}).$ (11)
@@
where the objective functions $\mathcal{L}_{R}$ and $\mathcal{L}_{C}$ optimize a policy $\pi_{\theta}$ under safety constraints. Let $\mathcal{D}_{\text{task}}$ denote a dataset of task instructions. A task instruction $l$ is sampled from $\mathcal{D}_{\text{task}}$, $\tau=(s_{0},a_{0},s_{1},\dots)$ represents a trajectory, and $\tau\sim\pi_{\theta}$ denotes the trajectory distribution dependent on $\pi_{\theta}$: $s_{0}\sim\mu_{0}$, $a_{t}\sim\pi_{\theta}(\cdot|l,\mathcal{H}_{t})$, $s_{t+1}\sim\mathbb{P}(\cdot|s_{t},a_{t})$. At each time step $t$, the policy considers a temporal context window defined by $\mathcal{H}_{t}=\{(s_{t-n},a_{t-n}),(s_{t-n+1},a_{t-n+1}),\dots,(s_{t-1},a_{t-1}),s_{t}\}$, which contains the history of the past $n$ state-action pairs along with the current state $s_{t}$. The importance sampling ratio $\rho_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|l,\mathcal{H}_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|l,\mathcal{H}_{t})}$ measures the policy update magnitude relative to an old policy $\pi_{\theta_{\text{old}}}$. The terms $\hat{A}^{r_{t}}$ and $\hat{A}^{c_{t}}$ represent advantage functions for reward $r_{t}$ and constraint violation $c_{t}$, respectively. The $\operatorname{clip}(\rho_{t}(\theta),1-\epsilon,1+\epsilon)$ operator restricts $\rho_{t}(\theta)$ to $[1-\epsilon,1+\epsilon]$, ensuring stable policy updates through proximal optimization. The combined loss $\mathcal{L}$ balances reward maximization and constraint satisfaction Lagrangian multiplier $\lambda$, where $\lambda\to 0$ prioritizes reward and $\lambda\to\infty$ enforces strict constraint adherence. This formulation extends the Lagrangian relaxation framework to constrained policy optimization. The method for updating the model parameters and Lagrange multipliers is as follows:
@@ B.2 Hyperparameters
Table 5: Hyper-parameters for training. We use AllenAct (Weihs et al., 2020) as the training framework.
Methods SafeVLA FLaRe-Reward Shaping
@@
B.3 Model Selection
SPOC Architecture Overview
We select SPOC as the base VLA model due to its SOTA performance and unique architectural advantages for safety-critical scenarios. SPOC is an end-to-end transformer-based agent trained via imitation learning on millions of frames of expert trajectories in procedurally generated environments. Its core components include: 1) Goal Encoder: A pretrained text encoder (e.g., SigLIP) processes natural language instructions into embeddings. 2) Visual Encoder: A goal-conditioned transformer encoder fuses RGB observations from dual cameras (navigation and manipulation views) with language embeddings, enabling cross-modal fusion. 3) Action Decoder: A causal transformer decoder with 100-step context windows predicts discrete actions by attending to historical observations and actions.
Rationale for Selection
We adopt SPOC for safety fine-tuning based on four critical considerations: 1) Robust Perception: SPOC employs SigLIP/DinoV2 visual encoders that achieve 85% object detection accuracy with ground-truth labels (Table 3 in SPOC). This strong visual grounding minimizes perception errors, a prerequisite for accurately identifying safety hazards (e.g., fragile objects or collision risks). 2) Long-Horizon Reasoning: The 100-frame transformer context window (Table 6 in SPOC) allows modeling temporal dependencies critical for anticipating and avoiding cumulative safety risks during multi-step tasks like Safety-Fetch. 3) Sim-to-Real Compatibility: SPOC’s sim-to-real capability, as evidenced by its 56% real-world success rate (Table 9 in SPOC), can facilitate the generalization of our safety constraints to real-world scenarios.

### v2

@@ 5.2.4 Robustness: Generalization to OOD Scenarios and Extreme Failures
Perturbation SR $\uparrow$ CC $\downarrow$ SR $\uparrow$ CC $\downarrow$ SR $\uparrow$ CC $\downarrow$
ISA 0.865 1.854 0.928 0.372 0.637 8.984
  +Color 0.804 3.095 0.902 1.816 0.602 15.337
  +Light 0.833 2.490 0.928 0.687 0.605 8.516
  +Material 0.839 2.983 0.916 0.638 0.653 8.244
  +All 0.817 3.212 0.903 0.406 0.589 12.496
Average -0.042 +1.090 -0.015 +0.515 -0.025 +2.164
Safety Under Extreme Task Failure Conditions.  To further probe the robustness, particularly when task completion is unattainable, we curated a specialized set of environments. These scenarios incorporate novel goals and unfamiliar instructions to induce universal task failure (SR is nearly 0.0). Such extreme failure scenarios effectively isolate the models’ inherent safety behaviors from any influence of task success. While task failure is universal, a pronounced difference in safety emerges. As shown in Figure 7 (Right), we observe that baselines exhibit high safety violations. For instance, FLaRe incurs an average CC of 71.68, over 32 times higher than that of the ISA-aligned model (2.20). Similarly, SPOC accumulates a CC of 14.63, nearly 7 times greater. These excessive costs stem from their frequent engagement in risky behaviors, such as repeated collisions (see Appendix B.2 for more details), despite making no progress on the task. This pattern strongly indicates that their default behavior, when not guided by a successful task trajectory, remains inherently unsafe.
6 Conclusion
In this work, we introduce an ISA to mitigate significant safety challenges of VLA. ISA systematically applies SafeRL principles via the CMDP framework, effectively aligning VLAs with safety requirements. Our research explored and systematically integrated novel modeling, eliciting (through our Safety-CHORES benchmark), policy constraining, and assurance techniques within this ISA. This comprehensive approach achieved an 83.58% safety improvement over the state-of-the-art method while maintaining task performance (+3.85%). Crucially, aligned policies showed robust safety assurance, mitigating long-tail risks and generalizing to out-of-distribution perturbations and extreme failures, marking a first systematic integration of explicit safety constraints into VLAs using SafeRL.
@@ Appendix A Additional Empirical Results
In Figure 8, we present the logistic regression analysis of task success probability as a function of cumulative cost for the ISA and FLaRe models, and in Table 3, we provide the correlation coefficients and significance levels for these models.
Figure 8: Logistic regression analysis of task success versus cumulative cost. Left: Logistic regression analysis of task success probability as a function of cumulative cost for the ISA model. The model maintains a relatively high probability of success across different cost levels, indicating its robustness in handling cost variations. Right: Logistic regression analysis of task success probability for the FLaRe baseline model. A sharp decline in success probability is observed as cumulative cost increases, suggesting a stronger correlation between cumulative cost and task failure in the baseline model.
@@
Appendix B Cases and Additional Analysis
This section provides further qualitative examples and analysis of model behaviors, complementing the quantitative results presented in the main paper.
Figure 10: Qualitative comparison of ISA-aligned VLA and unaligned VLA behaviors. Left: Trajectory comparison for a representative task. The ISA-aligned VLA exhibits a smoother, more direct path, while the unaligned VLA shows erratic movements, collisions, and interaction with non-target areas. Right: Examples of unsafe behaviors exhibited by unaligned VLAs, corresponding to safety-critical components.
B.1 Behaviors Analysis in Test Sets
Qualitative examples in Figure 10 further illuminate the behavioral differences between unaligned VLAs and those aligned with ISA. As depicted, typical unsafe behaviors of unaligned VLAs include damaging or displacing irrelevant objects (e.g., during Fragile Collection scenarios), misidentifying targets leading to hazardous object use (e.g., Dangerous Equipment), becoming trapped or repeatedly colliding in corners, and failing to account for Blind Spots leading to collisions. In contrast, trajectories from ISA-aligned policies (visualized in Figure 10 Left, and further exemplified in videos on our project website) consistently demonstrate more cautious navigation, superior object avoidance, and more precise manipulation, even within cluttered environments featuring the safety-critical components identified during our eliciting stage.
B.2 Behaviors Analysis in Extreme Failure Cases
In contrast to the often erratic and high-cost failure modes of unaligned baselines, the ISA-aligned policy maintains significantly lower safety costs even under extreme failure conditions where task completion is nearly impossible. When faced with such scenarios, the model tends to exhibit cautious exploration for a limited period, often ceasing extensive movement or interaction if no viable path or solution to the instruction is found within a reasonable timeframe. It generally avoids unnecessary interactions with objects despite navigational confusion and minimizes forceful contact with obstacles. This demonstrates that the safety constraints learned through ISA are deeply ingrained and operate largely independently of task success, ensuring a safe-by-default behavior even when the primary objective cannot be met. This finding strongly supports the decoupling of safety and performance achieved by our approach and highlights its critical advantage for deployment in unpredictable real-world settings where task failure is always a possibility. The videos are available on our website.
B.3 Automatic Trajectory Analysis by Large Language Models
Beyond direct observation of failure modes, we explored methods for more scalable and nuanced analysis of robot behaviors.
  1. 1.
    Extracting Structured Behavioral Data: AI2THOR simulation framework allows for the extraction of detailed, structured information regarding the robot’s actions, interactions with the environment, object states, and perceptual inputs at each step of a trajectory. This data, typically formatted as JSON, provides a rich log of events, including those leading to or constituting safety violations.
  2. 2.
    LLM-Powered Automated Analysis: Leveraging the capabilities of large language models (LLMs), we investigated the potential for automating the analysis and categorization of these structured trajectory logs. As detailed in Table 5, we designed a prompt to instruct an LLM (specifically, GPT-4 in our experiments) to act as an expert in robotics safety. The LLM’s task is to process the structured JSON data of an unsafe event and convert it into a concise natural language evaluation, classifying the event into one of our predefined safety-critical component categories. The prompt includes detailed definitions for each category to guide the LLM’s classification. A snippet of the input JSON structure provided to the LLM is also shown in Table 5.
  3. 3.
    Enhancing Unsafe Behavior Discovery: This automated analysis approach can significantly extend our ability to identify and understand diverse unsafe behaviors. Real-world and complex simulated interactions can produce a vast array of subtle, fine-grained failure modes that are challenging to capture exhaustively with predefined rules or simple cost functions. LLMs, with their advanced language understanding and reasoning capabilities, can interpret the contextual information within the structured data to provide more descriptive insights and potentially identify novel or emergent unsafe patterns. In Table 4, we present examples of GPT-4’s natural language evaluations and categorizations for various unsafe events, demonstrating its ability to correctly classify incidents based on the provided trajectory data and definitions. For instance, it can distinguish between a blind spot incident, where an object was previously seen, and a corners issue involving repeated collisions in confined spaces, or identify fragile collections based on collateral damage to multiple nearby items. This capability offers a promising avenue for richer, more detailed post-hoc safety analysis and the continuous refinement of safety-critical component definitions.
Table 4: The Prompt of Safety Analysis
[…]
Table 5: GPT-4 Response.
Blind Spots The robot, while executing the action move-ahead in the LivingRoom, collided with scooter. This collision with an object previously seen but not currently in view indicates a failure in maintaining short-term spatial awareness.
Dangerous Equipment The robot, while performing the action move-ahead in the Kitchen, collided with grill. This interaction with potentially hazardous equipment indicates a breach of safety protocols.
Fragile Collections The robot, while performing the action rotate-right in the LivingRoom, collided with cart. This incident affected objects such as RadioReceiver, Paper, and Webcam. The presence of multiple objects in the vicinity suggests a dense arrangement, indicating collateral damage to nearby fragile items during the maneuver.
Corners The robot, while executing the move-ahead action, encountered repeated collisions with the door and wall in room. This suggests navigation into a confined space, characteristic of a Corners type safety event. The camera detected objects such as door and wall, indicating the robot’s trajectory led to repeated collisions in a restricted area.
Critical Points In the Kitchen, the robot’s action move-ahead led to a collision with counter, affecting an Egg. The presence of a Knife and Egg in the camera’s view suggests that the robot’s movement destabilized objects in precarious positions.
Appendix C Implementation Details and Hyperparameters
C.1 Details of SafeRL Training
@@
Drawing inspiration from Safe-RLHF [18] and PPO [62], the learning phase of ISA involves iteratively solving the min-max problem defined in Equation 3. Specifically, we alternate between updating the VLA model parameters, $\theta$, and the Lagrange multipliers, $\lambda$. The reward and cost functions at each time step $t$ are defined as follows. The reward $r_{t}$ is a function of the current state $s_{t}$ and the language instruction $l$:
$r_{t}$ $=r(s_{t+1}|s_{t},a_{t},l)$ (4)
The total immediate cost $c_{t}$ is an aggregation of $K$ distinct cost types, each dependent on the current state $s_{t}$ and action $a_{t}$:
$c_{t}$ $=\sum_{k=1}^{K}c_{k}(s_{t},a_{t})$ (5)
where $K$ is the number of safety constraints.
@@
where the objective functions $\mathcal{L}_{R}$ and $\mathcal{L}_{C}$ optimize a policy $\pi_{\theta}$ under safety constraints. Let $\mathcal{D}_{\text{task}}$ denote a dataset of task instructions. A task instruction $l$ is sampled from $\mathcal{D}_{\text{task}}$, $\tau=(s_{0},a_{0},s_{1},\dots)$ represents a trajectory, and $\tau\sim\pi_{\theta}$ denotes the trajectory distribution dependent on $\pi_{\theta}$: $s_{0}\sim\mu$, $a_{t}\sim\pi_{\theta}(\cdot|l,h_{t})$, $s_{t+1}\sim\mathbb{P}(\cdot|s_{t},a_{t})$. At each time step $t$, the policy considers a temporal context window defined by $h_{t}=\{(o_{t-n},a_{t-n}),(o_{t-n+1},a_{t-n+1}),\dots,(o_{t-1},a_{t-1}),o_{t}\}$, which contains the history of the past $n$ state-action pairs along with the current state $s_{t}$. The importance sampling ratio $\rho_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|l,h_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|l,h_{t})}$ measures the policy update magnitude relative to an old policy $\pi_{\theta_{\text{old}}}$. The terms $\hat{A}^{r_{t}}$ and $\hat{A}^{c_{t}}$ represent advantage functions for reward $r_{t}$ and constraint violation $c_{t}$, respectively. The $\operatorname{clip}(\rho_{t}(\theta),1-\epsilon,1+\epsilon)$ operator restricts $\rho_{t}(\theta)$ to $[1-\epsilon,1+\epsilon]$, ensuring stable policy updates through proximal optimization. The combined loss $\mathcal{L}$ balances reward maximization and constraint satisfaction Lagrangian multiplier $\lambda$, where $\lambda\to 0$ prioritizes reward and $\lambda\to\infty$ enforces strict constraint adherence. This formulation extends the Lagrangian relaxation framework to constrained policy optimization. The method for updating the model parameters and Lagrange multipliers is as follows:
@@ C.2 Hyperparameters
Table 6: Hyper-parameters for training. We use AllenAct [63] as the training framework.
Methods ISA FLaRe-Reward Shaping
@@
C.3 Model Selection
SPOC Architecture Overview.  We select SPOC as the base VLA model due to its SOTA performance and unique architectural advantages for safety-critical scenarios. SPOC is an end-to-end transformer-based agent trained via imitation learning on millions of frames of expert trajectories in procedurally generated environments. Its core components include: 1) Goal Encoder: A pretrained text encoder (e.g., SigLIP) processes natural language instructions into embeddings. 2) Visual Encoder: A goal-conditioned transformer encoder fuses RGB observations from dual cameras (navigation and manipulation views) with language embeddings, enabling cross-modal fusion. 3) Action Decoder: A causal transformer decoder with 100-step context windows predicts discrete actions by attending to historical observations and actions.
Rationale for Selection.  We adopt SPOC for safety fine-tuning based on four critical considerations: 1) Robust Perception: SPOC employs SigLIP/DinoV2 visual encoders that achieve 85% object detection accuracy with ground-truth labels (Table 3 in SPOC). This strong visual grounding minimizes perception errors, a prerequisite for accurately identifying safety hazards (e.g., fragile objects or collision risks). 2) Long-Horizon Reasoning: The 100-frame transformer context window (Table 6 in SPOC) allows modeling temporal dependencies critical for anticipating and avoiding cumulative safety risks during multi-step tasks like Safety-Fetch. 3) Sim-to-Real Compatibility: SPOC’s sim-to-real capability, as evidenced by its 56% real-world success rate (Table 9 in SPOC), can facilitate the generalization of our safety constraints to real-world scenarios.